2. What is the pH of a solution when [H] = 8.99 x 1011 M?
Answer #1

solution [14+7= 8.99810-M. pH₂ -log[A+] = -log[8.99x6] - ll-log (8-997 = 11-0.9637 - 10.046 14 ~ 10.06